A trapzoid has height 8 feet and one base that measures 16 feet. If the area of the trapzoid is 168 ft, what is the length of the other base?
Answer by Boreal(15235)

A= (1/2)h (b1 +b2)
Looking for b2 (units are feet)
168 ft^2=(1/2) * 8 ft (b1+b2) ft units are fine.
Divide both sides by 4, since 8 *(1/2)=4
42 feet=(b1 + b2) feet
42=16+b2
Subtract 16 from both sides
b2=26 ft.
The median base is half of the sum or 21; 21*8=168 ft^2